The following information is available for Krane, Inc. Standard Actual DM Cost $6/foot $4,484 DM Quantity...

  1. The following information is available for Krane, Inc.

Standard

Actual

DM Cost

$6/foot

$4,484

DM Quantity

4 feet/unit

760 feet

DL Rate

$8.50/hour

$9.20/hour

DL Hours

2.4 hours/unit

450 hours

Variable OH Rate

$5.75/direct labor hour

$2,500

Variable OH Quantity

2.4 direct labor hours/unit

Fixed OH Cost (based on DLH)

$1,300 budgeted

$1,600

Fixed overhead is based on DLH. The capacity is 400 DLH. The company produced and sold 180 units of its product.

Calculate the following favorable or unfavorable:

  1. DM Cost Variance:                                                _________________________________
  2. DM Price Variance:                                               _________________________________
  3. DM Quantity Variance:                                        _________________________________
  4. DL Cost Variance:                                                  _________________________________
  5. DL Rate Variance:                                                  _________________________________
  6. DL Efficiency Variance:                                         _________________________________
  7. Variable OH Cost Variance:                                _________________________________
  8. Variable OH Spending Variance:                       _________________________________
  9. Variable OH Efficiency Variance:                      _________________________________
  10. Fixed OH Cost Variance:                                      _________________________________
  11. Fixed OH Budget Variance:                                 _________________________________
  12. Fixed OH Volume Variance:                                _________________________________
Answer #1

Solution:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
760 $5.90 $4,484.00 760 $6.00 $4,560.00 720 $6.00 $4,320.00
$76.00 F $240.00 U
Direct Material Price Variance Direct Material Qty variance
Direct material price variance $76.00 F
Direct material quantity variance $240.00 U
Direct material cost variance $164.00 U
Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
450 $9.20 $4,140.00 450 $8.50 $3,825.00 432 $8.50 $3,672.00
$315.00 U $153.00 U
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $315.00 U
Direct Labor Efficiency variance $153.00 U
Direct labor cost variance $468.00 U
Variable Overhead Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
450 $5.56 $2,500.00 450 $5.75 $2,587.50 432 $5.75 $2,484.00
$87.50 F $103.50 U
Variable overhead rate variance Variable overhead efficiency variance
Variable overhead rate variance $87.50 F
Variable overhead efficiency variance $103.50 U
Variable overhead cost variance $16.00 U
Fixed Overhead Cost Variance
Actual Fixed OH Cost Budgeted Fixed Overhead Standard Cost (FOH Applies)
SH* BR
$1,600.00 $1,300.00 432 $3.25 $1,404.00
$300.00 U $104.00 F
Fixed overhead Budget Variance Fixed overhead volume variance
Fixed overhead Budget Variance $300.00 U
Fixed overhead volume variance $104.00 F
Total Fixed overhead variance $196.00 F